OpenStudy (anonymous):

A projectile is fired straight upward with a velocity of 400 ft/sec. From physics, it's the distance above the ground after "t" seconds, given by s(t)=-16t^2 + 400t. a) What is the time elapsed when the projectile hits the ground? b) What is the "impact" velocity/ c) What is the maximum altitude achieved by the projectile? d) What is the acceleration at anytime "t"?

OpenStudy (turingtest):

we can use calculus, right?

OpenStudy (anonymous):

Yes

OpenStudy (turingtest):

first off what is s(t)=? when the projectile hits the ground

OpenStudy (turingtest):

?

OpenStudy (anonymous):

Well that was how the question was given to me...

OpenStudy (turingtest):

right, so I'm asking you: if s(t) represents the height of the ball, then what is s(t) when the ball hits the ground? Just use logic.

OpenStudy (anonymous):

0?

OpenStudy (turingtest):

sure! that's what "on the ground" means, right? so you need to solve\[s(t)=-16t^2+400t=0\] that is step one

OpenStudy (anonymous):

Okay cool so I get 25 sec

OpenStudy (anonymous):

For a

OpenStudy (turingtest):

right, now we need a formula for velocity hopefully you already know that v(t)=s'(t) so we need to take the derivative of s(t) and plug in t=25 into v(t) to get the impact velocity. Make sense?

OpenStudy (anonymous):

Yes, I understand, thank you, now I'm a little confused on c.

OpenStudy (anonymous):

But I understand d

OpenStudy (turingtest):

there are a few ways to see this... from a physics perspective, at the maximum height the velocity is zero, so we need to solve v(t)=0 from a calculus perspective we can use max/min techniques to find the maximum of s(t), which means setting s'(t)=0, which is the same thing. so physics and math are consistent, hooray!

OpenStudy (turingtest):

back in 5min if you have more questions...

OpenStudy (anonymous):

Okay thank you so much.

OpenStudy (turingtest):

ok a little preoccupied, I have to go again but do you have any quick questions?

OpenStudy (anonymous):

Um can you help differentiate y=2^3x

OpenStudy (turingtest):

this is a tolatally different problem, yes?

OpenStudy (turingtest):

totally*

OpenStudy (anonymous):

Yes, haha. sorry, you can go if your busy

OpenStudy (turingtest):

yeah five me another five or ten if you mean \[2^{3x}\]try to apply\[{d \over dx}(a^x)=a^x \ln a\]plus the chain rule \[f[(g(x)]'=f'[g(x)]g'(x)\] and see what you get. good luck!

OpenStudy (anonymous):

Thank you!